### v2

@@ 4.2.3 Development Trend Analysis
In addition, we also counted the publication time of the sample. We found that the number of publications has been fluctuating at a low value from 1999 to 2013, which may be related to the low level of information technology development at that time. Since 2017, with the rapid improvement of information technology, the number of published articles has also begun to increase significantly, and the blowout growth will be formed in 2023, reaching 30 articles. However, the number of published documents in 2020-2022 has dropped slightly, which we suspect may be related to the impact of the global novel coronavirus epidemic, and the development of relevant technologies is also limited. It is worth noting that the number of 2024 publications shown in the figure is less than 5, because we used the data up to January 2024 when compiling the data, so the 2024 data shown in the figure is not very valuable for reference. However, according to our trend forecast, the publication volume in the relevant field is likely to maintain significant growth in 2024 (see Figure 4).
Figure 4: Chart of Trends in the Number of Articles Published in the Field
@@ 5.2 Challenges and Limitations in Immersive Cultural Heritage
Secondly, sustainability issues are first reflected in terms of duration, technology and financial investment, which may be applied to the wider field of cultural heritage display [217]. In addition, users are unstable, changing participation, diverse cultural and political differences, and personalized dynamic needs mean that the digital application of cultural heritage needs more long-term and stable support. Third, the audience’s sense of experience will be affected by information overload, and the huge amount of information will make some experience groups feel at a loss. What is more obvious is that the full application of digitalization will lead to the limitation of cultural authenticity and physical experience, and it is difficult to fully convey the cultural background and deep meaning of some traditional crafts, resulting in defects in audience understanding [218]. Maintaining a balance between virtual content and the real world environment can be a challenge, and too many virtual elements can also distract visitors [219]. Fourth, security issues are reflected in the protection of cultural heritage and copyright issues, user personal data privacy and security issues [220]. In addition, for some users, the VR experience may cause motion sickness effects and reduce the quality of the experience [221].
@@ 5.3 Gaps and Risks in Current Immersive Cultural Heritage Practices
We also found some areas that have not been fully explored in the research, and these missing parts may have an important impact on the digital display and dissemination of cultural heritage. It is manifested in the misinterpretation of cultural heritage, the influence of purity in media experience, and the potential destructiveness of cultural heritage.
5.3.1 Overinterpretation or Misinterpretation of Cultural Heritage in Immersive Representations
Digital immersion in the presentation of cultural heritage can sometimes lead to over-interpretation or misunderstanding, thus deviating from the original intention of technical intervention [223]. For example, digital artifacts have the sole purpose of deepening the understanding and appreciation of real artifacts. When the focus is no longer on content but on technology, it may not be easy to use and distract visitors from cultural heritage, while leading to technological bias [224]. In addition, 3D reconstructions may overly rely on imagination or virtual additions, deviating from historical accuracy [225, 226]. Spatiotemporal discrepancies, contextual detachment of real data, and technical errors in digital databases can further distort public understanding and compromise authenticity [225].
@@ 6 Conclusion and Future Work
In this paper we have provided a comprehensive review of the application of immersive technologies in cultural heritage, identifying key opportunities, challenges, and gaps in the field. Analyzing over 5,000 articles, we distilled insights from 177 key papers to map the current landscape of immersive technologies, including VR, AR, and MR, and their use in cultural heritage. Our findings highlighted the transformative impact of these technologies in enhancing preservation, education, and cultural dissemination, offering personalized, interactive experiences that deepen audience engagement and promote cross-cultural dialogue. Immersive technologies enable innovative and accessible ways to present cultural heritage, expanding cultural appreciation and knowledge [220].
However, significant challenges remained, including technical barriers such as platform dependence and interoperability, sustainability concerns, and risks of diminishing cultural authenticity through digital reconstructions. Ethical concerns, such as data privacy and misrepresentation of cultural heritage, complicate the responsible use of these technologies. Misinterpretations through inaccurate reconstructions and over-reliance on technology, which can detract from the core media experience, further highlight the need for caution.

### v3

@@ 4.2.2 Development Trend Analysis
In addition, we also counted the publication time of the sample. Considering that digitization and immersive presentation in the field of cultural heritage have undergone significant development over time—moving from initial exploratory phases to more mature, diverse applications—we have opted for a longer timeframe to capture this evolution. We aim to cover the early emergence of related research, the iterative technological advancements, the accumulation of practical applications, and the most recent rapid developments in the field. We found that the number of publications has been fluctuating at a low value from 1999 to 2013, which may be related to the low level of information technology development at that time. Since 2017, with the rapid improvement of information technology, the number of published articles has also begun to increase significantly, and the blowout growth will be formed in 2023, reaching 30 articles. However, the number of published documents in 2020-2022 has dropped slightly, which we suspect may be related to the impact of the global novel coronavirus epidemic, and the development of relevant technologies is also limited. It is worth noting that the number of 2024 publications shown in the figure is less than 5, because we used the data up to January 2024 when compiling the data, so the 2024 data shown in the figure is not very valuable for reference. However, according to our trend forecast, the publication volume in the relevant field is likely to maintain significant growth in 2024 (see Figure 3).
Figure 3: Chart of Trends in the Number of Articles Published in the Field (1999-2024)
@@ 5.2 Challenges and Limitations in Immersive Cultural Heritage
Secondly, sustainability issues are first reflected in terms of duration, technology and financial investment, which may be applied to the wider field of cultural heritage display [217]. In addition, users are unstable, changing participation, diverse cultural and political differences, and personalized dynamic needs mean that the digital application of cultural heritage needs more long-term and stable support.
Third, the audience’s sense of experience will be affected by information overload, and the huge amount of information will make some experience groups feel at a loss. What is more obvious is that the full application of digitalization will lead to the limitation of cultural authenticity and physical experience, and it is difficult to fully convey the cultural background and deep meaning of some traditional crafts, resulting in defects in audience understanding [218]. Maintaining a balance between virtual content and the real world environment can be a challenge, and too many virtual elements can also distract visitors [219].
Fourth, security issues are reflected in the protection of cultural heritage and copyright issues, user personal data privacy and security issues [220]. In addition, for some users, the VR experience may cause motion sickness effects and reduce the quality of the experience [221].
@@ 5.3 Gaps and Risks in Current Immersive Cultural Heritage Practices
We also found some areas that have not been fully explored in the research, and these missing parts may have an important impact on the digital display and dissemination of cultural heritage. It is manifested in the misinterpretation of cultural heritage, the influence of purity in media experience, and the potential damage of cultural heritage.
5.3.1 Overinterpretation or Misinterpretation
of Cultural Heritage in Immersive Representations
Digital immersion in the presentation of cultural heritage can sometimes lead to over-interpretation or misunderstanding, thus deviating from the original intention of technical intervention [223]. For example, digital artifacts have the sole purpose of deepening the understanding and appreciation of real artifacts. When the focus is no longer on content but on technology, it may not be easy to use and distract visitors from cultural heritage, while leading to technological bias [224].
In addition, 3D reconstruction techniques have been implemented in many significant research projects for digital cultural heritage [225, 226]; however, these techniques sometimes rely excessively on imagination and/or visual additions [227, 228]. Spatiotemporal discrepancies, contextual detachment of real data, and technical errors in digital databases can further distort public understanding and compromise authenticity [225].
@@ 6 Conclusion and Future Work
In this paper we have provided a comprehensive review of the application of immersive technologies in cultural heritage, identifying key opportunities, challenges, and gaps in the field. Analyzing over 5,000 articles, we distilled insights from 177 key papers to map the current landscape of immersive technologies, including VR, AR, and MR, and their use in cultural heritage. Our findings highlighted the transformative impact of these technologies in enhancing preservation, education, and cultural dissemination, offering personalized, interactive experiences that deepen audience engagement and promote cross-cultural dialogue.
However, grand challenges remained, including technical barriers such as platform dependence and interoperability, sustainability concerns, and risks of diminishing cultural authenticity through digital reconstructions. Data privacy and misrepresentation of cultural heritage are critical considerations that must be addressed alongside other ethical issues. Ensuring responsible use of these technologies requires a balanced approach that respects both the technical and ethical dimensions. Furthermore, addressing these areas necessitates ongoing research to ensure that the development of digital and immersive display technologies honors the essence of cultural heritage and aligns with societal expectations.